OpenStudy (anonymous):

Judah and 4 of his friends shared some apples. Each of the five friends had 1 1/8 of an apple. How many apples did they eat?

OpenStudy (anonymous):

1 1/8 times 5

OpenStudy (anonymous):

9/8*5/1=45/8=5 and 5/8